OpenStudy (anonymous):

An object is launched at 20 m/s from a height of 65 m. The equation for the height (h) in terms of time (t) is given by h(t) = -4.9t2 + 20t + 65. What is the object's maximum height?

OpenStudy (zale101):

You can find the maximum height of the object from the vertex. Do you remember how to find the vertex of an equation? 

OpenStudy (anonymous):

i really don't know i'm sorry i'm so bad at math

OpenStudy (zale101):

You can either complete the square or find the vertex by this formula \(\Large Vertex=(h,k)=(\LARGE-\frac{b}{2a},c-\frac{b^2}{4a})\) Where ax^2+bx+c=0 is the standard form -4.9t^2 + 20t + 65=0 a=-4.9 b=20 c=65